SELF- AND H_{2}-BROADENING AND SHIFT COEFFICIENTS IN THE 2-0 BAND OF ^{12}C^{16}O

\underline{M. A. H. SMITH}, C. P. RINSLAND, Atmospheric Sciences, NASA Langley Research Center, Mail Stop 401A, Hampton, VA 23681-2199; A. W. MANTZ, Department of Physics, Astronomy and Geophysics, Connecticut College, New London, CT 06320; D. CHRIS BENNER and V. MALATHY DEVI, Department of Physics, The College of William and Mary, Box 8795, Williamsburg, VA 23187-8795.

Self- and hydrogen-broadening and pressure-shift coefficients for the first overtone band transitions of ^{12}C^{16}O at room temperature have been determined through analysis of nine high-resolution (0.005 cm^{−1}) absorption spectra. These spectra were recorded using the 1-m Fourier transform spectrometer (FTS) at the McMath-Pierce facility of the National Solar Observatory on Kitt Peak, Arizona. Eight of the nine spectra (including the three H_{2}-broadened spectra) were recorded using a high-purity CO sample having natural isotopic composition, and a single self-broadened spectrum was obtained using a CO sample having 99.999 atom % ^{12}C. A Pyrex cell with 10 cm path length was used to record all of the spectra. Self-broadened sample pressures ranged from about 3 to 507 Torr, and the H_{2}-broadened spectra had total pressures of 155 to 475 Torr. Because of the short path length of the cell, the volume mixing ratios of CO in H_{2} were relatively high, approximately 18% to 22%, to achieve measurable absorption in the 2-0 band. These mixing ratios necessitated the simultaneous determination of the CO self-broadening and self-shift coefficients along with the H_{2}-broadening and H_{2}-induced shift coefficients. We have determined these coefficients at room temperature, along with line positions and intensities, for the P(26) through R(22) ^{12}C^{16}O 2-0 transitions by fitting the entire spectral interval from 4130 to 4360 cm^{−1} in all nine spectra simultaneously. Our multispectrum nonlinear least-squares procedure^{a} was used to retrieve these spectral line parameters. The results will be compared to previous published values where possible.

aD. Chris Benner, C. P. Rinsland, V. Malathy Devi, M. A. H. Smith and D. Atkins, JQSRT \underline{53}, 705-721 (1995).